Grade 11Physical Chemistry

1 g of metal nitrate gave 0.86 g of metal sulphate. What is the equivalent weight of metal?

Manthan Asher
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago

Let a – nuclear weight of metal , 
N = valence of metal 
Weight of metal nitrate/Eq weight of metal + eq wt of metal nitrate = weight of sulfate/Eq wt of metal + eq wt of metal sulfate 
1/E + 62 = 0.86/E + 48 
E + 48 = 0.86 E + 53.32 
0.14 E = 5.32 
E = 38 .